OpenStudy (anonymous):

how many sides does a regular polygon have if each interior angle is 144

OpenStudy (anonymous):

Break up the polygon in your head. Maybe draw a picture of the polygon as a bunch of isosceles triangles that intersect at the center of the circle. If n stands for the number of sides on the polygon, the angle that goes out from the center is 360/n. That means that the other two angles added together are the angle of the polygon. And since they make up the angle of a triangle, they add to 180 degrees. So 120+360/n=180 360/n=180-120=60 60*n=360 n=360/60=6 hexagon So 140+360/n=180 360/n=180-140=40 40*n=360 n=360/40=9 nonagon So 144+360/n=180 360/n=180-144=36 36*n=360 n=360/36=10 decagon

OpenStudy (anonymous):

You can give medals by clicking 'Best Response'.